# appendixReferences

### v3

@@ A.2.1 Proof of Theorems 1, 2, and 3
A.2.2 Proof of Theorem 4
We prove (44), (45), and (46) separately, as follows:
@@ A.2.2 Proof of Theorem 4
[…]
To evaluate $\mathbb{E}a$, we observe that for $a\in\{0,1\}$, $\phi_{i}^{*}=1$, and $\tilde{f}_{ki}(T_{i})$ given by (65), it holds that
$\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}Y_{ki}\right]$ $=\mathbb{E}\left[\mathbb{E}\left\{\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}Y_{ki}|T_{i}\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{Y_{ki}|T_{i},\bar{Z}_{i}=a\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{\tilde{f}_{ki}(T_{i})|T_{i},\bar{Z}_{i}=a\right\}\right]$ (77)
@@
where (77) hold since $Y_{ki}=\tilde{f}_{ki}(T_{i})$ if $\phi_{i}^{*}=1$ and $X\in R_{i}$, which is implied by $\bar{Z}_{i}\in\{0,1\}$. It follows that $\mathbb{E}a$ and $\mathbb{E}b$ equal
$\mathbb{E}a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[\tilde{f}_{ki}(T_{i})]\\ \dfrac{1}{N}\displaystyle\sum_{i:\phi_{i}^{*}=1}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})]\end{array}\right],$ $\mathbb{E}b=\left[\begin{array}[]{l}1\\ 1\end{array}\right]$ (82)
and that $\hat{\Phi}_{k}^{\textup{Haj}}$ and $\Phi$ are given by
$\hat{\Phi}_{k}^{\textup{Haj}}$ $=N\left(\frac{a_{1}}{b_{1}}+\frac{a_{2}}{b_{2}}\right),$ $\Phi$ $=N\left(\frac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}+\frac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\right)$
@@
$\frac{1}{N}\left(\hat{\Phi}_{k}^{\textup{Haj}}-\Phi\right)$ $=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
@@
so that $N^{-1}(\hat{\Phi}_{k}^{\textup{Haj}}-\Phi)=h(\epsilon,b)$. Letting $\xi=(\epsilon,b)$, by mean value theorem it holds for some $\tilde{\xi}$ in the convex hull of $\xi$ and $\mathbb{E}\xi$ that
@@
As $\xi$ is an affine transformation of $\bar{u}$ defined in Lemma 1, it follows that
$\|\xi-\mathbb{E}\xi\|_{\infty}=O_{P}(N^{-1/2})$ (89)
@@
$\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)=O_{P}(N^{-1/2})$ (90)
@@
As $\tilde{\xi}$ is in the convex hull of $\xi$ and $\mathbb{E}\xi$, (89) implies that $\|\tilde{\xi}-\mathbb{E}\xi\|=O_{P}(N^{-1/2})$. By continuity of $\nabla h$ at $\mathbb{E}\xi$, it follows that $\|\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\|=O_{P}(N^{-1/2})$ as well, and hence that
$\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)=O_{P}(N^{-1})$ (91)
Substituting (90) and (91) into (87) proves that $N^{-1}(\hat{\Phi}_{k}-\Phi)=O_{P}(N^{-1/2})$, which is equivalent to the claim of (44) made by the lemma. ∎

### v4

@@ D.1 Proof of Propositions 1 - 5
D.2 Proof of Theorems 4 and 5
Proof of Theorem 4.
Given design neighborhoods $\Gamma_{i}$ satisfying (37) and (38), for $i\in[N]$ let $G_{i}=\{i\}$ and $\eta_{i}=\Gamma_{i}$. The conditions (50) - (56) of Assumption 2 can be verified as follows:
  1. 1.
    (50) and (53) follow from the definitions of $\{G_{i}\}$ and $\{\eta_{i}\}$
  2. 2.
    (51) and (52) follow because $\eta_{i}=\Gamma_{i}$, which satisfies (37) and (38).
  3. 3.
    (54) holds because it is assumed as (40)
  4. 4.
    To show (55) and (56), we observe that $|\Gamma_{i}\cup\Gamma_{j}|\leq 2d$, and hence by (39) and independence of the treatment assignments it holds that
    $P(X_{\Gamma_{i}\cup\Gamma_{j}}=x)\geq\min(p,1-p)^{2d}\qquad\forall x\in\{0,1\}^{|\Gamma_{i}\cup\Gamma_{j}|}$
    As a result, for any $(t_{i},t_{j})\in\operatorname{supp}(T_{i},T_{j})$ it holds that
    $P(T_{i}=t_{i},T_{j}=t_{j})\geq\min(p,1-p)^{2d}$ (83)
    where we have used that $(T_{i},T_{j})$ depend on $X$ only through $X_{\Gamma_{i}\cup\Gamma_{j}}$. Similarly, for any $(a,b,t_{i},t_{j})\in\operatorname{supp}(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ it also holds that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=t_{j})\geq\min(p,1-p)^{2d},$ (84)
    where we have used that $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depends on $X$ only through $X_{\Gamma_{i}\cup\Gamma_{j}}$. By (83) and (84), it follows that (55) and (56) hold for $\rho=\min(p,1-p)^{2d}$.
We prove Theorem 5 in two parts. First we prove the result under the original conditions A1 - A5. We then show that the result continues to hold when Conditions A3 and A4 are replaced by Condition A6, as stated by the theorem.
Proof of Theorem 5, using Conditions A1 - A5.
The conditions (50) - (56) of Assumption 2 can be verified as follows:
  1. 1.
    (50), (51), (52), and (54) hold because they are explicitly assumed by the conditions of the theorem.
  2. 2.
    (53) holds because $\bar{Z}_{i}$ and $T_{i}$ depend only on $X_{i}$, $X_{\mathcal{N}_{i}}$, and $\theta_{\eta_{g(i)}}$, all of which depend only on the treatment of unit $i$’s own and neighboring groups.
  3. 3.
    To show (55) and (56), we will show that for $\rho$ given by
    $\rho=p^{2d}\left(\frac{p}{1+(1-p)2d}\right)^{2d}$
    it holds that
    $P(T_{i}=t_{i},T_{j}=t_{j})$ $\geq\rho\qquad\forall(t_{i},t_{j})\in\operatorname{supp}(T_{i},T_{j})$ (85)
    which implies (55), and it also holds that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=t_{j})$ $\geq\rho\qquad\forall(a,b,t_{i},t_{j})\in\operatorname{supp}(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ (86)
    which implies (56).
To complete the proof, we now show (86), which also implies (85). Let the sets $S\subset[N]$ and $Q\subset[M]$ be given by
$S=\mathcal{N}_{i}\cup\mathcal{N}_{j}\qquad\textup{and}\qquad Q=\eta_{g(i)}\cup\eta_{g(j)}$
Because $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depend on $X$ and $\theta$ only through $X_{S}$ and $\theta_{Q}$, to show (86) it suffices to show that
$P(X_{S}=x,\theta_{Q}=\xi)$ $\geq\rho\qquad\forall(x,\xi)\in\operatorname{supp}(X_{S},\theta_{Q})$ (87)
To show (87), let $X_{S,i}$ denote the $i$th entry of $X_{S}$, let $X_{S,1:i-1}$ denote the first (i1)(i-1) entries of $X_{S}$, let $\bar{p}=\min(p,1-p)$, and let $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ denote terms to be defined shortly. It then holds for any $(x,\xi)\in\operatorname{supp}(X_{S},\theta_{Q})$ that
$P(X_{S}=x,\theta_{Q}=\xi)$ $\geq p^{2d}P(X_{S}=x|\theta_{Q}=\xi)$ (88)
$=p^{2d}\prod_{i=1}^{|S|}P(X_{S,i}=x_{i}|\theta_{Q}=\xi,X_{S,1:i-1}=x_{1:i-1})$ (89)
$=p^{2d}\prod_{i=1}^{|S|}\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ (90)
$\geq p^{2d}\left(\frac{\bar{p}}{1+(1-\bar{p})2d}\right)^{2d}$ (91)
by the following steps:
  1. 1.
    (88) holds because $\theta_{1},\ldots,\theta_{M}$ are indpendent and satisfy (58)
  2. 2.
    (89) writes $P(X_{S}|\theta_{Q})$ as a product of conditional probabilities
  3. 3.
    (90) holds because the entries of $X_{S}$ conditional on $\theta_{Q}$ are generated by sampling without replacement from each group, with $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ defined as follows. Let $\{\ell_{1},\ldots,\ell_{|S|}\}\subset[N]$ enumerate the units in $S$, so that $X_{S}=(X_{\ell_{1}},\ldots,X_{\ell_{|S|}})$ and $X_{S,i}=X_{\ell_{i}}$. Then it holds that
    $P(X_{S,i}=x_{i}|\theta_{Q}=\xi,X_{S,1:i-1}=x_{1:i-1})=\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$
    for $(a_{i},b_{i},k_{i},\tilde{k}_{i})$ given by
    $a_{i}$ $=\begin{cases}\theta_{g(\ell_{i})}&\textup{if }x_{i}=1\\ |G_{g(\ell_{i})}|-\theta_{g}(\ell_{i})&\textup{if }x_{i}=0,\end{cases}$ $b_{i}$ $=|G_{g(\ell_{i})}|$
    $k_{i}$ $=\sum_{j=1}^{i-1}1\{x_{j}=x_{i},g(\ell_{j})=g(\ell_{i})\},$ $\tilde{k}_{i}$ $=\sum_{j=1}^{i-1}1\{g(\ell_{j})=g(\ell_{i})\}$ (92)
    because the conditional probability of unit $\ell_{i}$ receiving treatment is equal to the probability of drawing a colored ball from an urn with $|G_{g(\ell_{i})}|$ balls, of which $\theta_{g(\ell_{i})}$ are colored, conditional on previously drawing without replacement a certain number of colored and non-colored balls corresponding to the treatment assignment of the units in $\{\ell_{1},\ldots,\ell_{i-1}\}$ who are in the same group as $\ell_{i}$.
  4. 4.
    To show (91) holds we observe that
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}=1\qquad\textup{ if }\theta_{g(\ell_{i})}=0\textup{ or }\theta_{g(\ell_{i})}=|G_{g(\ell_{i})}|$ (93)
    which holds because the treatment of unit $\ell_{i}$ can only take one value when all units in $\ell_{i}$’s group receive the same treatment, and
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}\geq\frac{\bar{p}}{1+(1-\bar{p})d}\qquad\textup{ if }\theta_{g(\ell_{i})}\neq 0\textup{ and }\theta_{g(\ell_{i})}\neq|G_{g(\ell_{i})}|$ (94)
    which holds by the following steps for $\theta_{g(\ell_{i})}\notin\{0,|G_{g(\ell_{i})}|\}$:
    $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ $\geq\frac{a_{i}-k_{i}}{b_{i}-k_{i}}$ (95)
    $\geq\frac{a_{i}-k_{i}}{\frac{a_{i}}{\bar{p}}-k_{i}}$ (96)
    $=\frac{\bar{p}(a_{i}-k_{i})}{(a_{i}-k_{i})+(1-\bar{p})k_{i}}$ (97)
    $\geq\frac{\bar{p}}{1+(1-\bar{p})2d}$ (98)
    where (95) because $\tilde{k}_{i}\geq k_{i}$ by their definitions in (92); (96) holds because $b_{i}=|G_{g(\ell_{i})}|\leq a_{i}/\bar{p}$ by (57); (97) follows from algebraic manipulation; and (98) holds because $k_{i}\leq|S|\leq 2d$, and because the expression
    $\frac{\bar{p}(a_{i}-k_{i})}{(a_{i}-k_{i})+(1-\bar{p})2d}$
    is increasing in $(a_{i}-k_{i})$, where $(a_{i}-k_{i})\geq 1$ since $x$ is in the support of $X_{S}$ conditional on $\theta_{Q}$. Plugging into (90) the bounds for $\frac{a_{i}-k_{i}}{b_{i}-\tilde{k}_{i}}$ given by (93) and (94) and bounding $|S|\leq 2d$ establishes (91).
Proof of Theorem 5, using Condition A6.
To verify the conditions (50) - (56) of Assumption 2 when Conditions A3 and A4 are replaced by Condition A6, we observe that
  1. 1.
    As before, (50), (51), (52), and (54) hold because they are explicitly assumed by the conditions of the theorem.
  2. 2.
    (53) holds because $\bar{Z}_{i}$ and $T_{i}$ depend only on $\theta_{\eta_{g(i)}}$, the number of treated units in $i$’s own and neighboring groups.
  3. 3.
    (55) holds because $T_{i}$ and $T_{j}$ are independent for $i\neq j$, under the conditions of the theorem.
  4. 4.
    To show (56), we observe that $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ depend on $X$ only through the group-level variables $\{\theta_{m}:m\in\eta_{g(i)}\cup\eta_{g(j)}\}$. Since $\{\theta_{m}\}_{m=1}^{M}$ are independent variables which satisfy the support condition (58), and $|\eta_{g(i)}\cup\eta_{g(j)}|\leq 2d$ by (51), it follows for any $(a,b,t_{i},t_{j})$ in the support of $(\bar{Z}_{i},\bar{Z}_{j},T_{i},T_{j})$ that
    $P(\bar{Z}_{i}=a,\bar{Z}_{j}=b,T_{i}=t_{i},T_{j}=T_{j})\geq p^{2d}$ (99)
    which implies (56) for $\rho=p^{2d}$.
D.3 Proof of Theorem 6
Preliminaries
it holds that
Lemma 1 applies McDiarmid’s inequality to show concentration of the terms appearing in $\hat{\tau}_{k}^{\textup{Haj}}$ for $k=1,2$.
Lemma 1.
Let Assumption 2 hold, and let $\phi^{*}\in\{0,1\}^{N}$ encode $\mathcal{I}$} by
$\phi_{i}^{*}=1\{i\in{\mathcal{I}}\},\qquad i\in[N]$ (102)
For $i\in[N]$ and $\bar{Z}_{i}$ given by (29), let $u_{i}$ denote the vector
and let $\bar{u}=\frac{1}{N}\sum_{i=1}^{N}u_{i}$. Then $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}\left(\frac{N}{\sqrt{M}}\right)$.
Proof.
Let $U_{m}$ for $m\in[M]$ be given by
$U_{m}=\frac{1}{N}\sum_{i\in G_{m}}u_{i}$
[…]
$P(|\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}|\geq\epsilon)\leq\exp\left(-\frac{2\epsilon^{2}}{M\left(\frac{dC}{M\rho}\right)^{2}}\right)$
and hence $\bar{u}_{\ell}-\mathbb{E}\bar{u}_{\ell}=O_{P}(M^{-1/2})$. Since $\bar{u}$ is 4-dimensional, it follows that $\|\bar{u}-\mathbb{E}\bar{u}\|_{\infty}=O_{P}(M^{-1/2})$ as well. ∎
Proof of Theorem 6
The claims of Theorem 6, namely (64) and (65), are proven separately as follows:
Proof of Theorem 6, Eq. (64).
It holds that
$\hat{\tau}_{k}^{\textup{Haj}}$ $=N-\left(\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}+\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{ki})\right)$
$\geq N-\left(\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}+\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot(1-Y_{ki})\right)$ (103)
$=N-\left(\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\cdot\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}\right)-N$ (104)
$=\begin{cases}-\Delta^{{\textup{Haj}}}&\textup{ if }k=1\\ \Delta^{{\textup{Haj}}}&\textup{ if }k=2\end{cases}$ (105)
where (103) holds because the terms in the summand are non-negative; (104) uses
$\sum_{i=1}^{N}\frac{N}{\hat{N}_{a}}\cdot\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})}=N,\qquad\forall\ a\in\{0,1\},$
and (105) uses the definition of $\Delta^{\textup{Haj}}$ given by (33). It follows that
$\max(\hat{\tau}_{1}^{\textup{Haj}},\hat{\tau}_{2}^{\textup{Haj}})$ $\geq\max(\Delta^{\textup{Haj}},-\Delta^{\textup{Haj}})$
$=|\Delta^{\textup{Haj}}|$
proving (64).
@@ D.3 Proof of Theorem 6
$a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\cdot Y_{ki}\phi_{i}^{*}\\ \\ \dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\cdot(1-Y_{ki})\phi_{i}^{*}\end{array}\right]$  $\qquad\textup{ and }\qquad b$ $=\left[\begin{array}[]{c}\dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}\\ \\ \dfrac{1}{N}\displaystyle\sum_{i=1}^{N}\dfrac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\end{array}\right]$ (112)
To evaluate $\mathbb{E}a$, we observe that for $z\in\{0,1\}$, $i\in{\mathcal{I}}$, and $\tilde{f}_{ki}(T_{i})$ given by (76), it holds that
$\mathbb{E}\left[\frac{1\{\bar{Z}_{i}=z\}}{P(\bar{Z}_{i}=z|T_{i})}Y_{ki}\right]$ $=\mathbb{E}\left[\mathbb{E}\left\{\frac{1\{\bar{Z}_{i}=z\}}{P(\bar{Z}_{i}=z|T_{i})}Y_{ki}|T_{i}\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{Y_{ki}|T_{i},\bar{Z}_{i}=z\right\}\right]$
$=\mathbb{E}\left[\mathbb{E}\left\{\tilde{f}_{ki}(T_{i})|T_{i},\bar{Z}_{i}=z\right\}\right]$ (113)
@@
where (113) hold since $Y_{ki}=\tilde{f}_{ki}(T_{i})$ if $i\in{\mathcal{I}}$ and $X\in R_{i}$, which is implied by $\bar{Z}_{i}\in\{0,1\}$. It follows that $\mathbb{E}a$ and $\mathbb{E}b$ equal
$\mathbb{E}a$ $=\left[\begin{array}[]{l}\dfrac{1}{N}\displaystyle\sum_{i\in{\mathcal{I}}}\mathbb{E}[\tilde{f}_{ki}(T_{i})]\\ \dfrac{1}{N}\displaystyle\sum_{i\in{\mathcal{I}}}\mathbb{E}[1-\tilde{f}_{ki}(T_{i})]\end{array}\right],$ $\mathbb{E}b=\left[\begin{array}[]{l}1\\ 1\end{array}\right]$ (118)
and that $\hat{\tau}_{k}^{\textup{Haj}}$ and $\tau$ are given by
$\hat{\tau}_{k}^{\textup{Haj}}$ $=N\left(\frac{a_{1}}{b_{1}}+\frac{a_{2}}{b_{2}}\right),$ $\tau$ $=N\left(\frac{\mathbb{E}a_{1}}{\mathbb{E}b_{1}}+\frac{\mathbb{E}a_{2}}{\mathbb{E}b_{2}}\right)$
@@
$\frac{1}{N}\left(\hat{\tau}_{k}^{\textup{Haj}}-\tau\right)$ $=\frac{\epsilon_{1}}{b_{1}}+\frac{\epsilon_{2}}{b_{2}}$
@@
so that $N^{-1}(\hat{\tau}_{k}^{\textup{Haj}}-\tau)=h(\epsilon,b)$. Letting $\xi=(\epsilon,b)$, by mean value theorem it holds for some $\tilde{\xi}$ in the convex hull of $\xi$ and $\mathbb{E}\xi$ that
@@
As $\xi$ is an affine transformation of $\bar{u}$ defined in Lemma 1, it follows from that Lemma that
$\|\xi-\mathbb{E}\xi\|_{\infty}=O_{P}(M^{-1/2})$ (125)
@@
$\nabla h(\mathbb{E}\xi)^{T}(\xi-\mathbb{E}\xi)=O_{P}(M^{-1/2})$ (126)
@@
As $\tilde{\xi}$ is in the convex hull of $\xi$ and $\mathbb{E}\xi$, (125) implies that $\|\tilde{\xi}-\mathbb{E}\xi\|=O_{P}(M^{-1/2})$. By continuity of $\nabla h$ at $\mathbb{E}\xi$, it follows that $\|\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\|=O_{P}(M^{-1/2})$ as well, and hence that
$\left(\nabla h(\tilde{\xi})-\nabla h(\mathbb{E}\xi)\right)^{T}(\xi-\mathbb{E}\xi)=O_{P}(M^{-1})$ (127)
Substituting (126) and (127) into (123) proves that $N^{-1}(\hat{\tau}_{k}-\tau)=O_{P}(M^{-1/2})$, which is equivalent to the claim of (65) made by the lemma. ∎
D.4 Proof of Theorem 7
In our proof of Theorem 7, we will use $C_{ij}(T_{i},T_{j})$ to abbreviate the covariance estimates appearing in $\hat{V}_{k}$. It is given by
To prove the theorem, we will use Lemmas 2 - 7, whose proofs will be given following the proof of Theorem 7. Lemma 2 bounds the values of $\{C_{ij}\}$.
Lemma 2.
Let Assumption 2 hold and let $C_{ij}$ be given by (128). Let $\mathcal{J}_{0}$ and $\mathcal{J}_{1}$ denote the sets
$\mathcal{J}_{0}$ $=\{i,j\in{\mathcal{I}}:|\Gamma_{i}\cap\Gamma_{j}|=0\}$
$\mathcal{J}_{1}$ $=\{i,j\in{\mathcal{I}}:|\Gamma_{i}\cap\Gamma_{j}|\neq 0\}$
It holds w.p. 1 that
$\max_{i,j\in\mathcal{J}_{0}}C_{ij}(T_{i},T_{j})$ $\leq O\left(\frac{\log M}{M}\right)$ (129)
$\max_{i,j\in\mathcal{J}_{1}}C_{ij}(T_{i},T_{j})$ $\leq O(1)$ (130)
Lemmas 3 and 4 are used to prove (67) of Theorem 7.
Lemma 3.
Let $C_{ij}$ be given by (128). If $i,j\in{\mathcal{I}}$ and $P(\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}|T_{i},T_{j})=0$, then $C_{ij}(T_{i},T_{j})\leq 0$
Lemma 4.
Lemmas 5 and 6 bound terms appearing in our variance estimators, and will be used to prove Lemma 2.
Lemma 5.
Let Assumption 2 hold, and let $v_{i}$ for $i\in[N]$ denote
If $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|=0$, then
If $i,j\in{\mathcal{I}}$ and $|\Gamma_{i}\cap\Gamma_{j}|\neq 0$, then
$\mathbb{E}\left[v_{i},v_{j}|T_{i}T_{j}\right]$ $\leq\frac{1}{\rho^{2}}\mathbb{E}\left[\frac{N^{2}}{\hat{N}_{Y_{ki}}\hat{N}_{Y_{kj}}}|T_{i},T_{j},\bar{Z}_{i}=Y_{ki},\bar{Z}_{j}=Y_{kj}\right].$ (133)
Lemma 6.
Let Assumption 2 hold. Then
$\max_{i\in{\mathcal{I}}}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki}\right]$ $=1+O\left(\frac{\log M}{M}\right)$ (134)
$\max_{i,j\in{\mathcal{I}}}\mathbb{E}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=1+O\left(\frac{\log M}{M}\right)$ (135)
$\max_{i,j\in{\mathcal{I}}}{\operatorname{Var}\,}\left[\frac{N}{\hat{N}_{Y_{ki}}}|T_{i},\bar{Z}_{i}=Y_{ki},T_{j},\bar{Z}_{j}=Y_{kj}\right]$ $=O\left(\frac{1}{M}\right)$ (136)
Lemma 7 bounds the extent to which the design neighborhoods $\{\Gamma_{i}\}_{i=1}^{N}$ may overlap, and the number of design neighborhoods to which a group $G_{m}$ may belong.
Lemma 7.
Let Assumption 2 hold. It holds that
$\max_{i\in[N]}\sum_{j}1\{|\Gamma_{i}\cup\Gamma_{j}|\neq 0\}$ $\leq\frac{d^{2}CN}{M}$ (137)
$\max_{m\in[M]}\sum_{i}1\{G_{m}\subseteq\Gamma_{i}\}$ $\leq\frac{dCN}{M}$ (138)
Proof of Theorem 7
The claims of Theorem 7, namely (66), (67), and the conditions under which $\hat{V}_{k}$ is unbiased, are proven separately as follows.